OpenStudy (emendez10):

Find the angle between the two space curves below at their point of intersection. r1(t) = <3t-1, t^(2) +1, square root(t+2)> r2(t)=<2t^(2)+3, 1-4t, t+3>

OpenStudy (eliesaab):

Notice \[ \text{r2}(-1)=\text{r1}(2)=\{5,5,2\}\]

OpenStudy (eliesaab):

So (5,5,2) is where they intersect

OpenStudy (eliesaab):

Use the formula \[ \cos(\theta) = \frac{\vec{x}\cdot \vec{y}}{||\vec{x}|| \ ||\vec{y}||} \] where \[ \vec x=r1(2)\\ \vec y=r2(-1) \]

OpenStudy (irishboy123):

don't you want the angle between the tangent vectors?